OpenStudy (anonymous):

the slope of a line is 1/8. The point (-8, -5) is on the line. Determine the equation of this line.

OpenStudy (anonymous):

y = m(x-a) + b for (a,b) on the line with slope m so y = 1/8 (x-(-8)) -5 y = x/8 +1 -5 -> y = x/8 -4